OpenStudy (anonymous):

Find the product of (x + 7)2. x2 − 14x + 49 x2 + 14x + 49 x2 − 49 x2 + 49

OpenStudy (anonymous):

(x+7)^2 = (x+7)(x+7)= x(x+7) + 7(x+7) Now simplify

OpenStudy (anonymous):

\[(x + 7)^2\] apply the power of 2 to each factor: x^2 + 7^2 = \[x^2 + 49\]

OpenStudy (anonymous):

@jazy , check again. You are wrong

OpenStudy (anonymous):

whoops, I forgot to factor:/ Thanks for letting me know @shivam_bhalla !

OpenStudy (anonymous):

np :)